# Fine-Grained Domain Generalization with Feature Structuralization

### v2

@@ I Introduction
Figure 1: Instance of multi-granularity knowledge. Four animals are categorized into various classes across three granularity levels according to their commonalities and specificities.
In FGDG tasks, fine-grained categories exhibit multi-granularity structures built upon the commonalities and specificities between categories. This aspect is overlooked by previous DG methods. In contrast, humans exhibit a higher level of generalizability by leveraging these structured semantics for learning and recognizing. According to family resemblances [9], feature integration theory [10], and lexical structure [11] in cognitive psychology, when learning to recognize objects, people focus primarily on the salient targets and inductively disentangle the commonalities and specificities among them. Commonalities represent basic visual concepts, whereas specificities correspond to individual conceptual extensions. Both serve as the foundational elements for object classification and vary across distinct granularity levels [12, 13, 14]. For instance, as Fig. 1 shows, four animals can be reclassified into categories of Equidae and Bovidae based on their commonalities and specificities from a higher $c_{1}$ granularity level, respectively. At $c_{1}$ level, zebra and horse share certain characteristics, such as long necks, while manifesting specificities at fine-grained level $f_{0}$. It can be observed that the common and specific features, which constitute the intrinsic characteristics of objects, can be facilitated by incorporating multi-granularity knowledge. Understanding the commonalities and specificities between categories enhances the generalizability, as both focus more on the intrinsic features of the objects themselves.
@@
To formulate FS, we devise a pipeline consisting of two essential steps: disentanglement and alignment. We disentangle the three parts according to their channel indices. A decorrelation optimization function is employed to minimize the mutual information between them. Three similarity-based losses are then proposed for the alignment of common and specific parts with the aid of multi-granularity knowledge. Two of these are leveraged to pull commonalities closer across diverse granularities, while the third serves to create distance among the specific components. Furthermore, a Feature Structuralized Domain Generalization model (FSDG) is constructed by incorporating a prediction calibration method. Three model variations are expanded in conjunction with Convolutional Neural Network (CNN), Transformer, and Multi-Layer Perceptron (MLP).
@@ II-A Domain Generalization
Research on alleviating OOD issues has been conducted from various perspectives [2, 20]. DA, as a straightforward solution, directly utilizes target domain data as auxiliary knowledge to train the model [7]. However, collecting or even identifying target data before deploying the model is unbearable, and when applied to a third domain, the model still encounters failures [5]. Therefore, it is more appropriate to study the DG problem without training the target domain data. In recent years, numerous DG strategies have been proposed, such as learning domain-invariant representations [21, 22], data augmentation [23, 24], optimization strategies [25, 26, 27], and disentangled representations [28].
Robust domain-invariant representations emphasize computing the consistencies across multi-source domains [5]. Rosenfeld et al. [3] acquired domain-invariant features with the help of causal analysis. Literature [29] computed a kind of rationale invariance to enhance DG performance. Data augmentation methods can be used to generate multiple domain data for single-source DG problems [30]. [31] employed adversarial training to generate images with unseen styles. Zhou et al. [23] mixed the statistics of two images in the image or feature level to generate new styles. Wang et al. [25] improved DG capability by implicitly aligning the gradient directions between the empirical risk and the perturbed loss. Zhang et al. [1] introduced a training strategy called Lookahead, which involves weight interpolation, to explore flat minima for DG issues. In [32], two separate encoders were learned in an adversarial way to capture identity and domain information, respectively. Low-rank decomposition on weight matrices is applied in [33] to identify features that are more generalizable.
@@ II-B Fine-Grained Visual Categorization
Fine-grained visual categorization (FGVC) strives to discriminate subtle distinctions between subordinate categories within the same root category [6]. It presents a formidable challenge due to the inherent dilemma of minimal inter-class variation coupled with significant intra-class variation. Researchers endeavoring to handle this problem can be classified into three main paradigms: feature enhancement-based [34], localization-based [35], and external knowledge-based methods [36]. These methods mainly focus on enhancing the fine discriminative characteristics. For example, FET-FGVC [34] proposed a feature-enhanced Transformer, cooperating with Graph Neural Networks, to improve the performance of FGVC. Zheng et al. [35] firstly localized key object parts within images and then found discriminative clues on object parts. Literature [36] incorporated object structure knowledge into the transformer to upgrade discriminative representations. As for fine-grained generalization issues, PAN [7] harnessed multi-granularity knowledge, aligning the final multi-granular predictions to improve fine-grained DA performance. WDAN [37] utilized a weighted discriminative adversarial network to release the fine-grained DA issue. However, these methods neglect feature invariance between categories, and the single-source FGDG problem has not been explored.
@@ III-A Problem Statement
Domain Generalization: Let $\mathcal{X}=\left\{(x)\right\}$ be the input space and $\mathcal{Y}=\left\{(y)\right\}$ the label space. In the context of DG, $\mathcal{D_{S}}=\left\{\mathcal{S}^{s}=\left\{\left(x^{s},y^{s}\right)\right\}\right\}_{s=1}^{M}$ and $\mathcal{D_{T}}=\left\{\mathcal{T}\right\}$ denote $M$ source domains and an unseen target domain, respectively. We train a model $f:\mathcal{X}\rightarrow\mathcal{Y}$ utilizing only source domain data $\mathcal{D_{S}}$. The goal of DG is to minimize the prediction error of $f$ on the unseen target domain $\mathcal{D_{T}}$. Classical DG approaches work in multi-source scenarios (i.e., $M>1$), learning invariant representations of multiple source domains. In contrast, single-source DG (i.e., $M=1$) is a more realistic problem in open environments, as it is hard to ensure that invariant features remain consistent in the target domain, especially given the limited number of available source domains for training.
Fine-Grained Domain Generalization: In the FGDG problem, a model $f$ is trained on $M$ source domains $\mathcal{D_{S}}=\left\{\mathcal{S}^{s}=\left\{\left(x^{s},y_{f}^{s},\left.y_{g}^{s}\right|_{g=1}^{G-1}\right)\right\}\right\}_{s=1}^{M}$ and tested on a target domain $\mathcal{D_{T}}=\left\{\mathcal{T}\right\}$. The label space of $\mathcal{D_{S}}$ is constructed in a multi-granular format with a total $G$-layer hierarchy (i.e., $\mathcal{Y}=\left\{(y_{f},\left.y_{g}\right|_{g=1}^{G-1})\right\}$). Traditional DG can be seen as a particular kind of FGDG whose label space has only the fine-grained labels $y_{f}$ without the coarse-grained labels $\left.y_{g}\right|_{g=1}^{G-1}$, where $g$ is a notation of granularity level and subscript $f$ represents the fine-grained content (i.e., $g=0$). The objective of FGDG is to minimize the fine-grained prediction error on the unseen target domain $\mathcal{D_{T}}$. In this paper, we enhance FGDG performance in a single-source scenario (i.e., $M=1$), taking into account a more realistic situation in which multi-source fine-grained structured data is more difficult to fabricate. Additionally, single-source FGDG can demonstrate the efficacy of our method more effectively.
@@ III-B Disentanglement and Decorrelation of Three Partitions
$\left\{\mathcal{F}_{g}^{c},\mathcal{F}_{g}^{p},\mathcal{F}_{g}^{n}\right\}=Disentangle(\mathcal{F}_{g}),$ (1)
where $\mathcal{F}_{g}^{c}=\left\{\left.{f}_{g}^{c,i}\right|_{i=1}^{d^{c}}\right\}$ denotes $d^{c}$ common features, $\mathcal{F}_{g}^{p}=\left\{\left.{f}_{g}^{p,i}\right|_{i=1}^{d^{p}}\right\}$ reflects the specific segment, $\mathcal{F}_{g}^{n}=\left\{\left.{f}_{g}^{n,i}\right|_{i=1}^{d^{n}}\right\}$ is the confounding part, and $d=d^{c}+d^{p}+d^{n}$, respectively.
In this paper, we disentangle features at the conceptual level according to their channel indices, as the interpretability analysis shows that channels can express various concepts [18]. This procedure is also characterized by its conciseness, practicality, and computational efficiency. It allows us to explicitly identify the index positions of the functionality of the three parts, fortifying the transparency of the model.
@@
where $MEAN(X,dim)$ is the average operator of tensor $X$ along the $dim$ dimension, $F_{g}^{c}\in\mathbb{R}^{B\times d^{c}\times wh}$, $F_{g}^{p}$ and $F_{g}^{n}$ are slices of the stacked 3-d tensor $F_{g}\in\mathbb{R}^{B\times d\times wh}$ along the channel dimension, with extraction ratios set to $r^{c}$, $r^{p}$ and $r^{n}$. The resulted $P_{g}^{c}\in\mathbb{R}^{B\times wh}$, $P_{g}^{p}$ and $P_{g}^{n}$ are then stacked as $P_{g}^{all}\in\mathbb{R}^{B\times 3\times wh}$. The objective loss function of decorrelation among three segments is defined as
$\mathcal{L}_{dec}=\frac{1}{B}\frac{1}{G}\sum_{b=0}^{B-1}\sum_{g=0}^{G-1}SUM({\left\langle P_{b,g}^{all},P_{b,g}^{all}\right\rangle\over\|P_{b,g}^{all}\|_{2}\|P_{b,g}^{all}\|_{2}}-\mathbf{I}),$ (3)
where $SUM(X)$ is the summation operator among all elements of tensor $X$, and $\mathbf{I}$ is a identity matrix. Cosine similarity is utilized as the metric because it can directly quantify orthogonality. Other metrics, such as Hilbert-Schmidt independence criterion (HSIC) [39] and Euclidean distance [40], are also tested, and results are presented in Section IV.
@@ III-C Alignment for Commonality, Specificity and Prediction
In particular, the output features of $\left.T_{g}\right|_{g=0}^{G-1}$ are initially stacked as a 4-d tensor $F\in\mathbb{R}^{G\times B\times d\times wh}$. All features are tokenized, resulting in $wh$ tokens in dimension $d$, which enables transformer-based variants of FS. The common slice is sequentially extracted from $F$ based on Eq. (1). The first constraint, denoted as $\mathcal{L}_{cs}$, measures how similar the common features of the same samples are across each pair of neighboring granularities. This distance is formulated as
$S_{cs}=\frac{1}{B}\frac{1}{G-1}\sum_{b=0}^{B-1}\sum_{g=0}^{G-2}SUM({\left\langle{f}_{b,g}^{c},{f}_{b,g+1}^{c}\right\rangle\over\|{f}_{b,g}^{c}\|_{2}\|{f}_{b,g+1}^{c}\|_{2}}),$ (4)
where ${f}_{b,g}^{c}\in\mathbb{R}^{d^{c}\times wh}$ belongs to ${g}_{th}$ granularity of ${b}_{th}$ image.
For the second constraint, we locate sub-granularity samples of every parent class, progressing from the highest parent granularity to the lowest. Given a batch of images $x$ in $\left.m_{g}\right|_{g=0}^{G-1}$ classes, the samples belonging to $q_{th}$ parent category $y_{g+1}^{q}$ are of different classes at sub-granularity $g$, denoted by $\left.x^{q,k}_{g}\right|_{k=1}^{K}$, where $k$ is the sub-category, and $K=m_{g}^{q}$ is the according number of sub-categories. For every sub-category, we compute the GAP of the obtained feature tensor $F_{B_{k,q,g}}^{c}\in\mathbb{R}^{B_{k,q,g}\times d^{c}\times wh}$ to represent the fundamental common centroid:
$P_{k,q,g}^{c}=MEAN(GAP(F_{B_{k,q,g}}^{c}),dim=0),$ (5)
where $B_{k,q,g}$ depicts samples that belong to $q_{th}$ category at granularity $g+1$ and $k_{th}$ sub-category at $g$. $P_{k,q,g}^{c}$ represents the centroid of the common characteristics of samples belonging to the current sub-granularity category. The second constraint, denoted as $\mathcal{L}_{cd}$, forces the sub-centroids of a given parent category to move closer. The total distance among the stacked sub-centroid tensors $P_{q,g}^{c}\in\mathbb{R}^{K\times d^{c}}$ for all parent categories and all granularities is formulated as
$S_{cd}=\frac{1}{G-1}\frac{1}{Q_{g}}\frac{1}{K}\sum_{g=0}^{G-2}\sum_{q=0}^{Q_{g}-1}SUM({\left\langle P_{q,g}^{c},P_{q,g}^{c}\right\rangle\over\|P_{q,g}^{c}\|_{2}\|P_{q,g}^{c}\|_{2}}-\mathbf{I}),$ (6)
where $Q_{g}$ is the number of parent categories at granularity $g$.
Specificity Alignment regulates the second group of features to match the distinctive characteristics between objects, denoted as $S_{p}$. The main idea is that the specific representations of different categories at the same granularity should be maximally distinct. We compute the specificity centroid as
@@
The total distance among the stacked centroids $P_{g}^{p}\in\mathbb{R}^{K\times d^{p}}$ for all granularities can be formulated as
$S_{p}=\frac{1}{G-1}\frac{1}{K}\sum_{g=0}^{G-1}SUM({\left\langle P_{g}^{p},P_{g}^{p}\right\rangle\over\|P_{g}^{p}\|_{2}\|P_{g}^{p}\|_{2}}-\mathbf{I}).$ (8)

### v3

@@ I Introduction
Figure 1: Instance of multi-granularity knowledge. Four animals are categorized into various classes across three granularity levels according to their commonalities and specificities, as described in the green region. The numbers represent species labels: labels 4, 5, 6, and 7 correspond to the $f_{0}$ granularity level; 2 and 3 to the $c_{1}$ level; and 1 to the coarsest $c_{2}$ level. For commonalities, the number outside parentheses indicates the parent category, while the two inside are its sub-categories.
@@
In FGDG tasks, fine-grained categories exhibit multi-granularity structures built upon the commonalities and specificities between categories. This aspect is overlooked by previous DG methods. In contrast, humans exhibit a higher level of generalizability by leveraging these structured semantics for learning and recognizing. According to information decomposition theory of the brain [9], family resemblances [10], feature integration theory [11], and lexical structure [12] in cognitive psychology, when learning to recognize objects, people focus primarily on the salient targets and inductively disentangle the commonalities and specificities among them. Commonalities represent basic visual concepts, whereas specificities correspond to individual conceptual extensions. Both serve as the foundational elements for object classification and vary across distinct granularity levels [13, 14, 15]. For instance, as Fig. 1 shows, four animals can be reclassified into categories of Equidae and Bovidae based on their commonalities and specificities from a higher $c_{1}$ granularity level, respectively. At $c_{1}$ level, zebra and horse share certain characteristics, such as long necks, while manifesting specificities at fine-grained level $f_{0}$. It can be observed that the common and specific features, which constitute the intrinsic characteristics of objects, can be facilitated by incorporating multi-granularity knowledge. Understanding the commonalities and specificities between categories enhances the generalizability, as both focus more on the intrinsic features of the objects themselves.
@@
To formulate FS, we devise a pipeline consisting of two essential steps: disentanglement and alignment. We disentangle the three parts according to their channel indices. A decorrelation optimization function is employed to minimize the mutual information between them. It enables the confounding segment to learn complementary and redundant information for recognition [21, 22, 23]. Three similarity-based losses are then proposed for the alignment of common and specific parts with the aid of multi-granularity knowledge. Two of these are leveraged to pull commonalities closer across diverse granularities, while the third serves to create distance among the specific components. Furthermore, a Feature Structuralized Domain Generalization model (FSDG) is constructed by incorporating a prediction calibration method. Three model variations are expanded in conjunction with Convolutional Neural Network (CNN), Transformer, and Multi-Layer Perceptron (MLP).
@@ II-A Domain Generalization
Research on alleviating OOD issues has been conducted from various perspectives [2, 24]. DA, as a straightforward solution, directly utilizes target domain data as auxiliary knowledge to train the model [7]. However, collecting or even identifying target data before deploying the model is unbearable, and when applied to a third domain, the model still encounters failures [5]. Therefore, it is more appropriate to study the DG problem without training the target domain data. In recent years, numerous DG strategies have been proposed, such as learning domain-invariant representations [25, 26], data augmentation [27, 28], optimization strategies [29, 30, 31], and disentangled representations [32, 33].
Robust domain-invariant representations emphasize computing the consistencies across multi-source domains [5]. Rosenfeld et al. [3] acquired domain-invariant features with the help of causal analysis. Literature [34] computed a kind of rationale invariance to enhance DG performance. Data augmentation methods can be used to generate multiple domain data for single-source DG problems [35]. Zhou et al. [27] mixed the statistics of two images in the image or feature level to generate new styles. Wang et al. [29] improved DG capability by implicitly aligning the gradient directions between the empirical risk and the perturbed loss. DualVAE [36] utilized a dual-path VAE architecture to disentangle latent factors and trained a classifier on top of the semantic factor for DG issues. Low-rank decomposition on weight matrices is applied in [37] to identify features that are more generalizable.
@@ II-B Fine-Grained Visual Categorization
Fine-grained visual categorization (FGVC) strives to discriminate subtle distinctions between subordinate categories within the same root category [6]. It presents a formidable challenge due to the inherent dilemma of minimal inter-class variation coupled with significant intra-class variation. Researchers endeavoring to handle this problem can be classified into three main paradigms: feature enhancement-based [38], localization-based [39, 40], and external knowledge-based methods [41]. These methods mainly focus on enhancing the fine discriminative characteristics. For example, FET-FGVC [38] proposed a feature-enhanced Transformer, cooperating with Graph Neural Networks, to improve the performance of FGVC. Zheng et al. [39] firstly localized key object parts within images and then found discriminative clues on object parts. Literature [41] incorporated object structure knowledge into the transformer to upgrade discriminative representations. As for fine-grained generalization issues, PAN [7] harnessed multi-granularity knowledge, aligning the final multi-granular predictions to improve fine-grained DA performance. WDAN [42] utilized a weighted discriminative adversarial network to release the fine-grained DA issue. Besides, LDRE [43] utilized VLMs to retrieve a target image based on the commonalities between the reference and candidates. However, these methods neglect feature invariance between categories, and the single-source FGDG problem has not been explored.
@@ III-A Problem Statement
Domain Generalization: Let $\mathcal{X}=\left\{(x)\right\}$ be the input space and $\mathcal{Y}=\left\{(y)\right\}$ the label space. In the context of DG, $\mathcal{D_{S}}=\left\{\left(\mathcal{X}^{s},\mathcal{Y}^{s}\right)\right\}_{s=1}^{M}$ and $\mathcal{D_{T}}=\left\{\mathcal{T}\right\}$ denote $M$ source domains and an unseen target domain, respectively. We train a model $f:\mathcal{X}\rightarrow\mathcal{Y}$ utilizing only source domain data $\mathcal{D_{S}}$. The goal of DG is to minimize the prediction error of $f$ on the unseen target domain $\mathcal{D_{T}}$. Classical DG approaches work in multi-source scenarios (i.e., $M>1$), learning invariant representations of multiple source domains. In contrast, single-source DG (i.e., $M=1$) is a more realistic problem in open environments, as it is hard to ensure that invariant features remain consistent in the target domain, especially given the limited number of available source domains for training.
Fine-Grained Domain Generalization: The label space of $\mathcal{D_{S}}$ is constructed in a multi-granular format with a total $G$-layer hierarchy (i.e., $\mathcal{Y}=\left\{(y_{f},\left.y_{g}\right|_{g=1}^{G-1})\right\}$). Traditional DG can be seen as a particular kind of FGDG whose label space has only the fine-grained labels $y_{f}$ without the coarse-grained labels $\left.y_{g}\right|_{g=1}^{G-1}$, where $g$ is a notation of granularity level and subscript $f$ represents the fine-grained content (i.e., $g=0$). The objective of FGDG is to minimize the fine-grained prediction error on the unseen target domain $\mathcal{D_{T}}$. In this paper, we enhance FGDG performance in a single-source scenario (i.e., $M=1$), taking into account a more realistic situation in which multi-source fine-grained structured data is more difficult to fabricate. Additionally, single-source FGDG can demonstrate the efficacy of our method more effectively.
@@ III-B Disentanglement and Decorrelation of Three Partitions
$\left\{\mathcal{F}_{g}^{c},\mathcal{F}_{g}^{p},\mathcal{F}_{g}^{n}\right\}=Disentangle(\mathcal{F}_{g}).$ (1)
In this paper, we disentangle features at the conceptual level according to their channel indices, as the interpretability analysis shows that channels can express various concepts [19]. This procedure is also characterized by its conciseness, practicality, and computational efficiency. It allows us to explicitly identify the index positions of the functionality of the three parts, fortifying the transparency of the model. Therefore, $\mathcal{F}_{g}^{c}=\left\{{f}_{g}^{1},\cdots,{f}_{g}^{{d^{c}}}\right\}$ has $d^{c}$ common features, $\mathcal{F}_{g}^{p}=\left\{{f}_{g}^{d^{c}+1},\cdots,{f}_{g}^{{d^{c}}+{d^{p}}}\right\}$ reflects the ${d^{p}}$ specific features, $\mathcal{F}_{g}^{n}=\left\{{f}_{g}^{d^{c}+{d^{p}}+1},\cdots,{f}_{g}^{{d}}\right\}$ is the confounding segment, where $d=d^{c}+d^{p}+d^{n}$, the superscript of feature $f$ denotes its index, and the subscript of $f$ indicates the granularity level to which it belongs, respectively.
@@
where $MEAN(X,dim)$ is the average operator of tensor $X$ along the $dim$ dimension, $F_{g}^{c}\in\mathbb{R}^{B\times d^{c}\times wh}$, $F_{g}^{p}\in\mathbb{R}^{B\times d^{p}\times wh}$ and $F_{g}^{n}\in\mathbb{R}^{B\times d^{n}\times wh}$ are slices of the stacked 3-d tensor $F_{g}\in\mathbb{R}^{B\times d\times wh}$ along the channel dimension, with extraction ratios set to $r^{c}$, $r^{p}$ and $r^{n}$. The resulted $P_{g}^{c}\in\mathbb{R}^{B\times wh}$, $P_{g}^{p}$ and $P_{g}^{n}$ are then stacked as $P_{g}^{all}\in\mathbb{R}^{B\times 3\times wh}$. To effectively harness the parallel computing capabilities of GPUs, we compute the covariance of $P_{b,g}^{all}\in\mathbb{R}^{3\times wh}$ of each sample $b$ in the batch $B$ as the distance between the 3 segments:
$\mathcal{L}_{dec}=AVG({\left\langle P_{b,g}^{all},P_{b,g}^{all}\right\rangle\over\|P_{b,g}^{all}\|_{2}\|P_{b,g}^{all}\|_{2}}-\mathbf{I}).$ (3)
Cosine similarity is utilized as the metric because it directly quantifies orthogonality. Other metrics, such as Hilbert-Schmidt independence criterion (HSIC) [45] and Euclidean distance [46], are also tested, and results are presented in Section IV. Considering that the prototype vectors exhibit unity correlation with themselves (the diagonal elements), we subtract an identity matrix $\mathbf{I}$. The elements of the covariance matrix, which subtracts self-correlation, yield the inter-correlation between the prototypes. The objective loss function of the decorrelation is averaged (i.e. the $AVG(\cdot)$ operator) among all samples and granularities (subscripts $b$ and $g$).
@@ III-C Alignment for Commonality, Specificity and Prediction
In particular, the first constraint, denoted as $\mathcal{L}_{cs}$, measures how similar the common features of the same samples are across each pair of neighboring granularities. This distance is formulated as
$S_{cs}=AVG({\left\langle{f}_{b,g}^{c},{f}_{b,g+1}^{c}\right\rangle\over\|{f}_{b,g}^{c}\|_{2}\|{f}_{b,g+1}^{c}\|_{2}}),$ (4)
where ${f}_{b,g}^{c}\in\mathbb{R}^{d^{c}\times wh}$ represents the common feature tensor belongs to ${g}_{th}$ granularity of ${b}_{th}$ image. This distance is averaged across all samples and all parent granularities (amounting to a total of $G-1$ levels).
For the second constraint, we locate sub-granularity samples of every parent class, progressing from the lowest parent granularity to the highest. Given a batch of images $x$, assume there are a total of $Q_{g+1}$ classes (parent classes) at parent granularity level $g+1$. For the $q_{th}$ parent class, it comprises a total of $K_{q,g}$ subcategories at the sub-granularity level $g$. For each subcategory $k$, we locate all samples belonging to this subcategory, which are denoted as $B_{k,q,g}$. Accordingly, we derive the feature tensor $F_{B_{k,q,g}}\in\mathbb{R}^{B_{k,q,g}\times d\times wh}$ for these samples. We then compute the GAP of the common segment $F_{B_{k,q,g}}^{c}\in\mathbb{R}^{B_{k,q,g}\times d^{c}\times wh}$ of the feature tensor to represent the fundamental common sub-centroid:
$P_{k,q,g}^{c}=MEAN(GAP(F_{B_{k,q,g}}^{c}),dim=0).$ (5)
$P_{k,q,g}^{c}\in\mathbb{R}^{d^{c}}$ represents the centroid of the common characteristics of the current sub-category $k$ belonging to the parent category $q$. The second constraint, denoted as $\mathcal{L}_{cd}$, forces the sub-centroids of a given parent category to move closer. Similar to Eq. 3, we compute the covariance matrix of the stacked sub-centroid tensor $P_{q,g}^{c}\in\mathbb{R}^{K_{q,g}\times d^{c}}$ and subtract the self-correlation of each sub-centroid:
$S_{cd}=AVG({\left\langle P_{q,g}^{c},P_{q,g}^{c}\right\rangle\over\|P_{q,g}^{c}\|_{2}\|P_{q,g}^{c}\|_{2}}-\mathbf{I}).$ (6)
The distance is also averaged across all parent categories and all granularities.
Specificity Alignment regulates the second group of features to match the distinctive characteristics between categories, denoted as $\mathcal{L}_{p}$. The main idea is that the specific representations of different categories at the same granularity should be maximally distinct, learning the differences between categories. Similarly, we locate all samples that belong to category $k$ at granularity $g$ and derive their specific feature tensor $F_{B_{k,g}}^{p}\in\mathbb{R}^{B_{k,g}\times d^{p}\times wh}$. The specificity centroid is defined as
@@
The averaged distance among the stacked centroids $P_{g}^{p}\in\mathbb{R}^{K_{g}\times d^{p}}$ for all granularities can be formulated as
$S_{p}=AVG({{\left\langle P_{g}^{p},P_{g}^{p}\right\rangle\over\|P_{g}^{p}\|_{2}\|P_{g}^{p}\|_{2}}-\mathbf{I}}).$ (8)
After these disentanglement, decorrelation, and alignment operations, the three partitions focus on recognizing different semantics. It is worth noting that no constraints are imposed on the confounding segment. This segment maintains the model’s redundancy and randomness, given the class-related nature of the first two partitions. The $\mathcal{L}_{dec}$ ensures that the third partition learns complementary information for recognition.
@@
Algorithm 1 Pseudocode of the FSDG Approach
1:Source-domain training samples $x^{s}$; Multi-granularity knowledge $\left(y_{f}^{s},\left.y_{g}^{s}\right|_{g=1}^{G-1}\right)$; coarse feature extractor $E_{c}$; fine feature extractor $E_{f}$; number of layers of the multi-granularity structure $G$; multi-granularity transition layers $\left.T_{g}\right|_{g=0}^{G-1}$; multi-granularity classifier $\left.C_{g}\right|_{g=0}^{G-1}$; channel dimensions $d$; extraction ratios $r^{c}$, $r^{p}$ and $r^{n}$.
2:Well-trained fine-branch networks $E_{f}^{\ast}$, $T_{0}$ and $C_{0}$;
3:for iteration = 1 : $N$ do
4:     Compute granular features $\left.\mathcal{F}_{g}\right|_{g=0}^{G-1}$ of $G$ transition layers;
5:     for g = 0 : $G-1$ do
6:         Disentangle $\mathcal{F}_{g}$ into 3 semantical segments $\mathcal{F}_{g}^{c}$, $\mathcal{F}_{g}^{p}$ and $\mathcal{F}_{g}^{n}$;
7:         Compute prototypes of the 3 segments according to Eq. (2);
8:         Compute distance among the 3 segments (in Eq. (3)); $\triangleright$ to get $\mathcal{L}_{dec}$
9:         if $g<G-1$ then
10:              Compute distance between ${f}_{b,g}^{c}$ and ${f}_{b,g+1}^{c}$ at every pair of adjacent granularities (in Eq. (4)); $\triangleright$ to get $S_{cs}$
11:              for each parent class $q$ at granularity $g+1$ do
12:                  for each sub-class $k$ at granularity $g$ that belongs to current parent class $q$ do
13:                       Find all samples that belong to current sub-class $k$ and parent class $q$;
14:                       Compute sub-centroid $P_{k,q,g}^{c}$ of the common segments of these samples according to Eq. (5);
15:                  end for
16:                  Compute distance between the sub-centroids $P_{q,g}^{c}$ of current parent class $q$ (in Eq. (6)); $\triangleright$ to get $S_{cd}$
17:              end for
18:         end if
19:         for each class $k$ at granularity $g$ do
20:              Find all samples that belong to class $k$;
21:              Compute the centroid $P_{k,g}^{p}$ of the specific segments of these samples according to Eq. (7);
22:         end for
23:         Compute distance between the centroids $P_{g}^{p}$ (in Eq. (8)); $\triangleright$ to get $S_{p}$
24:     end for
25:     Aggregate and average all the distances to derive $\mathcal{L}_{dec}$, $S_{cs}$, $S_{cd}$ and $S_{p}$, respectively;
26:     Optimize $\mathcal{L}_{lf}$ and $\mathcal{L}_{c}$ in Eq. (9) and Eq. (10);
27:     Optimize $\mathcal{L}_{FS}$ in Eq. (11) by minimizing $\mathcal{L}_{dec}$ and $S_{p}$, and maximizing $S_{cs}$ and $S_{cd}$;
28:end for
The pseudocode for the proposed FSDG approach is presented in Algorithm 1, where the core steps and formulas are elaborated in detail. It is worth noting that the aforementioned processes are parallelized to enhance GPU efficiency by applying the constraints on the covariance matrix.